a person takes part in the medical trial that tests the effect of medicine on a disease half the people are given medicine and the other half are given a suger pill which has no effect on disease the medicine has 60% chance of curing someone but people who do not get the medicine still have 10%chance getting cured determine the probality that the person will get well and assuming the person got cured determine the probability that it was not through the medication
Answer #1

1) P(get cured ) =P(treatment)*P(get well|treatment)+P(sugar pill)*P(get well|sugar pill)

=0.5*0.6+0.5*0.1 =0.35

2)

P(sugar pill(not through medication) given get cured )

=P(sugar pill)*P(get well|sugar pill)/P(get cured )

=0.5*0.1/0.35

=0.1429